GRAVITATIONAL CARNATION

Do you know little by little you are living inside me But with you there is a transfiguration^{∗} For instance we drink wine,^{†} as if a carnation is falling inside us A tree is working tick tock beside us My stomach was, my mind was—this little bit remains.^{‡}

You lean towards the carnation, I take and pass it to you voilà And you pass it more beautiful to another The other passes it along n’est-ce pas So goes the carnation from hand to hand.

You see then, you and I are growing a passion, I’m touching on you, I’m warming up to you, this is not that Look how, seven colors become as one We are silently uniting.

English interpretation^{§} by Ayşe and David from:

YERÇEKİMLİ KARANFİL

Biliyor musun az az yaşıyorsun içimde Oysaki seninle güzel olmak var Örneğin rakı içiyoruz, içimize bir karanfil düşüyor gibi Bir ağaç işliyor tıkır tıkır yanımızda Midemdi aklımdı şu kadarcık kalıyor.

Sen o karanfile eğilimlisin, alıp sana veriyorum işte Sen de bir başkasına verıyorsun daha güzel O başkası yok mu bir yanındakine veriyor Derken karanfil elden ele.

Görüyorsun ya bir sevdayı büyütüyoruz seninle Sana değiniyorum, sana ısınıyorum, bu o değil Bak nasıl, beyaza keser gibisine yedi renk Birleşiyoruz sessizce.

Edip Cansever, 1957

∗Literally growing beautiful.

†Literally rakı.

‡That is, This little bit of my stomach and mind has remained.

§For Annabelle and Ken, June, 2004.